Two balls are dropped to the ground from different heights.One is dropped 1.5 second after the other,but they both strike the ground at the same time,5 second after the first is dropped (a)What is the difference in the heights from which they are dropped. (b)From what height was the first ball dropped?

Grade:11

1 Answers

Arun
5 years ago
a)Let initially the first ball be at height h​1, and takes t1 time to reach  the ground = 5 second, and 
let initially the second ball be at height h​2, and takes t2 time to reach  the ground = 5-2= 3second
for first Ball                                                            For second ball
s = ut+1/2gt2                                                                          s = ut+1/2gt2  
h1 =  0 +1/2 g (5)2                                                h2 = 0 +1/2g(3)2
h1 =  25g/2   ...........1>                                          h2   = 9g/2      .............2>
now the difference in height is h1  - h2  =  ​ 25g/2   -  ​9g/2
                                                          = 16g/2  =8g 
                                                           8 x 9.8 =78.4 meter
  b)first ball dropped from the height = h1 = 25g/2 =122.5 m